OpenStudy (anonymous):

use the point slope form to write an equation of a line that passes through (3, -4) and is perpendicular to 2x-3y=5

OpenStudy (anonymous):

3x+2y=1

OpenStudy (anonymous):

find m of 2x-3y=5 (rearrange equation in standard form) y=(2/3)x-(5/3) so m=2/3

OpenStudy (anonymous):

solve the first equation in point slope form 2x-3y=5 where y = 2/3 x -5/3 Perpendicular slope you take the reciprocal and change the sign so your knew slope is m=-3/2

OpenStudy (anonymous):

perperndicular lines have reciprocal gradient to that of the given since m=-2/3 then new equation will be y=(2/3)x-5/3

OpenStudy (anonymous):

flip your slope remainder

OpenStudy (anonymous):

the sign that is